ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 178, 2010

Alaminos, J., J. Extremera and A. R. Villena Approximately zero-product-preserving maps

1

Baryshnikov, Y. and D. Romik Enumeration formulas for Young tableaux in a diagonal strip

157

Chong, C. T. and T. A. Slaman The Theory of the α degrees is undecidable

229

Extremera, J. See J. Alaminos Fox, J., J. Pach and C. D. Tóth Turán-type results for partial

orders and intersection graphs of convex sets

29

Humphries, S. P., B. L. Kerby and K. W. Johnson Fusions of character tables III: fusions of cyclic groups and a generalisation of a condition of Camina

325

Izmestiev, I. The Colin de Verdière number and graphs of polytopes  427
Jacquet, H. Distinction by the quasi-split unitary group  269

Johnson, K. W. See S. P. Humphries Katz, M. G. and S. Shnider Cayley form, comass, and triality iso-

morphisms

187

Kerby, B. L. See S. P. Humphries Levin, M. B. On low discrepancy sequences and low discrepancy er-

godic transformations of the multidimensional unit cube

61

Liu, J. and P. Sarnak Integral points on quadrics in three variables whose coordinates have few prime factors

393

Lovász, L. and B. Szegedy Testing properties of graphs and functions

113

Marigo, R. a Hartogs type theorem for separately CR functions  107

Mishchenko, S. and A. Valenti Varieties with at most quadratic growth

209

Pach, J. See J. Fox Podoski, K. and B. Szegedy On finite groups whose derived sub-

group has bounded rank

51

Romik, D. See Y. Baryshnikov Sarnak, P. See J. Liu Shnider, S. See M. G. Katz Slaman, T. A. See C. T. Chong Szegedy, B. See K. Podoski Szegedy, B. See L. Lovász Tóth, C. D. See J. Fox Valenti, A. See S. Mishchenko Vezzoni, L. Connections on contact manifolds and contact twistor

space

253

Villena, A. R. See J. Alaminos

Wu, J. The functor A^{min} for (p−1)-cell complexes and EHP sequences  349